OpenStudy (anonymous):

What is the derivative of e^(2x) I am confused because the derivative e^x is e^x. But they say you need to use the chain rule. Which is f'(g(x))*g'(x). I have used that on exponents but in this I just get, 2x * e^(2x-1) *e Which is not the answer =/

OpenStudy (anonymous):

2e^(2x)

OpenStudy (amistre64):

\[y=e^{2x}\] \[ln(y)=2x\] \[y'/y = 2\] \[y'=2y;\ but\ y=e^{2x}\] \[y' = 2e^{2x}\]

OpenStudy (anonymous):

So I just have to use logs, to differentiate it.

OpenStudy (shayaan_mustafa):

@amistre64 will you kindly elaborate elaborate your 3rd step i.e. y'/y=2 how did you get this? please.

OpenStudy (s3a):

(Chain) rule: d/dx e^[f(x)] = df(x)/dx * e^[f(x)] Examples: 1) d/dx e^x = dx/dx * e^x = 1 * e^x = e^x 2) d/dx e^(2x) = d(2x)/dx * e^(2x) = 2 * e^(2x)

OpenStudy (s3a):

Shayaan_Mustafa, if you treat y as a function (and not just a regular variable) then you take the derivative of y as if it's just a regular variable and you turn the y into a y'.

OpenStudy (s3a):

Example if y = y(x), d/dx (y^2) = 2y * y'

OpenStudy (anonymous):

1) d/dx e^x = dx/dx * e^x = 1 * e^x = e^x why does e^x = 1?

OpenStudy (anonymous):

o wait

OpenStudy (s3a):

d(x)/dx = 1

OpenStudy (s3a):

in this case f(x) = x.

OpenStudy (s3a):

This being e^x.

OpenStudy (anonymous):

ok , I kinda get it. So all I have to do is use the chain rule, why did he use the ln rule above?

OpenStudy (s3a):

He was proving it.

OpenStudy (s3a):

In other words, the very first time mathematicians had to do this, they had to justify why what I said works.

OpenStudy (s3a):

I think.

OpenStudy (s3a):

If someone can confirm, that'd be great (for me).

OpenStudy (anonymous):

ok lol =P

OpenStudy (s3a):

:)

OpenStudy (anonymous):

Thanks for the help everyone, I'm a little closer to understanding it.